• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 11:39:09 on February 23, 2009

    Entry number 262859

    This entry is a followup to: 262827

    keyword=re: BB Scint TDCs shifted (fixed via ROC5 power cycle)

    Power cycled ROC5 and the problem seems to have been resolved. (At least the cosmics TDC spectra look OK now.)